OpenStudy (jen543):

1 algebra question? will fan

OpenStudy (jen543):

(x+4)^2/(x-4) divided by (x^2-16)/(4x-16)

OpenStudy (3mar):

Could I help you?

OpenStudy (jen543):

yes

OpenStudy (3mar):

That is with my pleasure! Can you type your equation here?

OpenStudy (jen543):

\[(x+4)^2/(x-4) \div (x^2 - 16)/(4x-16)\]

OpenStudy (3mar):

Do you mean that? \[\LARGE \frac{ \frac{ (x+4)^2 }{ (x-4) } }{\frac{ (x^2-16) }{ (4x-16)}}\]

OpenStudy (jen543):

yeah but theres a division sign in between

OpenStudy (3mar):

That is represented by the main fraction between (x-4) and (x^2-16)

OpenStudy (3mar):

@Jen543 Do you follow?

OpenStudy (3mar):

Let me know where you are back!